OpenStudy (anonymous):

I WILL GIVE MEDAL AND BECOME A FAN 5x + 3y = -2 Isolate y: ______ on both sides ______ on both sides 5x + 3y =-2 3y = ______ y = ______ The rate of change is _____. Please explain this to me simply, thank you very much! :)

OpenStudy (mathmale):

Hey, Poco, you're looking for the SLOPE of this straight line. SLOPE can be interepreted as a "rate of change." In this case, were we to solve for y, the slope, m, would represent how much y changes if x changes by 1 unit. Can you complete the work: In other words, if you have 5x + 3y = -2, and you subtract 5x from both sides, what do you get? And how would you solve that result for y? Once you have that equation, how would you find (identify) the slope/rate of change?

OpenStudy (anonymous):

How do I subtract 5x from both sides? i

OpenStudy (mathmale):

5x + 3y = -2 -5x = -5x --------------- 3y = -2 - 5x. Now please solve for y by division: divide both sides of the equation by 3.